OpenStudy (anonymous):

3 = –(–y + 6) need help A. –9 B. –3 C. 3 D. 9

ganeshie8 (ganeshie8):

\(\large 3 = –(–y + 6)\) is same as \(\large 3 = –1(–y + 6)\)

ganeshie8 (ganeshie8):

distribute that -1 on right hand side

OpenStudy (anonymous):

im sorry i dont get you

ganeshie8 (ganeshie8):

distributive law of multiplication : \[\large a(b+c) = ab+ac\]

ganeshie8 (ganeshie8):

\[\large \begin{align}3&=-1(-y+6)\\~\\&=(-1)(-y) + (-1)(6)\\~\\&=y - 6\end{align}\]

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

\[\large 3 = y - 6\] add 6 both sides to isolate `y`

ganeshie8 (ganeshie8):

we're not done yet

ganeshie8 (ganeshie8):

what do you get when you add `6` to both sides ?

OpenStudy (anonymous):

8 and 6y?

ganeshie8 (ganeshie8):

3+6 = 9 right ?

ganeshie8 (ganeshie8):

\[\large 3 = y-6\] add 6 to both sides \[\large 3+6 = y-6+6\] \[\large 9 = y\]

ganeshie8 (ganeshie8):

so the value of unknown variable, y, is 9

OpenStudy (anonymous):

oh thank you so much can you help me with one more ?

ganeshie8 (ganeshie8):

I'll try... but are you sure you're able to follow all the mess that i wrote above ? :)